# Wi-CBR: WiFi-based Cross-domain Behavior Recognition via Multimodal Collaborative Awareness

### v1

@@ Appendix A More Implementation Details
Widar3.0  [7] is one of the most widely used datasets for WiFi-based gesture recognition. The dataset collection device consists of one transmitter and six receivers, each equipped with an Intel 5300 wireless card. Each receiver has three antennas arranged in a line. The operating frequency of the device is 5.825 GHz, with a sampling rate of 1000 packets per second. Widar3.0 mainly consists of two types of datasets. Firstly, there are 12,000 gesture samples commonly used in human-computer interaction (16 users × 5 locations × 5 orientations × 6 gestures × 5 instances). Secondly, there are 5,000 gesture samples representing digits 0-9 on a horizontal plane (2 users × 5 locations × 5 orientations × 10 gestures × 5 instances). To ensure a fair comparison, we only utilized 4,500 gesture samples (9 users × 5 locations × 5 orientations × 6 gestures × 5 instances) from the Widar3.0 dataset for source-free domain adaptation gesture recognition. The detailed descriptions of Widar3.0 dataset are provided in Tab. 4
Table 4: Detailed Descriptions of Widar3.0 Dataset
@@
1st (Classroom) 9 1: Push-Pull; 2: Sweep; 3: Clap; 4: Slide; 5: Draw-O(Horizontal); 6: Draw-Zigzag(Horizontal); 7: Draw-N(Horizontal); 8: Draw-Triangle(Horizontal); 9: Draw-Rectangle(Horizontal) 5 5 10,125
2nd (Hall) 4 1: Push-Pull; 2: Sweep; 3: Clap; 4: Slide; 5: Draw-O(Horizontal); 6: Draw-Zigzag(Horizontal) 5 5 3,000
3rd (Office) 4 1: Push-Pull; 2: Sweep; 3: Clap; 4: Slide; 5: Draw-O(Horizontal); 6: Draw-Zigzag(Horizontal) 5 5 3,000
Table 5: Detailed Descriptions of XRF55 Dataset
@@
Scene 1 30 1: Drawing circle; 2: Drawing cross; 3: Pushing; 4: Pulling; 5: Swiping left; 6: Swiping right; 7: Swiping up; 8: Swiping down Changeable Changeable 4,800
Scene 2 3 1: Drawing circle; 2: Drawing cross; 3: Pushing; 4: Pulling; 5: Swiping left; 6: Swiping right; 7: Swiping up; 8: Swiping down Changeable Changeable 480
Scene 3 3 1: Drawing circle; 2: Drawing cross; 3: Pushing; 4: Pulling; 5: Swiping left; 6: Swiping right; 7: Swiping up; 8: Swiping down Changeable Changeable 480
Scene 4 3 1: Drawing circle; 2: Drawing cross; 3: Pushing; 4: Pulling; 5: Swiping left; 6: Swiping right; 7: Swiping up; 8: Swiping down Changeable Changeable 480
XRF55  [56]is a large multi-radio frequency dataset forhuman action analysis, consisting of 42.9K samples of 55 action categories related to human-computer interaction collected from 39 users. To verify the effectiveness of the proposed Wi-CBR, we selected only the samples related to gesture actions collected using WiFi devices (e.g., drawing a circle, drawing a cross, pushing, pulling, swiping left, swiping right,swiping up, and swiping down) for the experiments. Among them, 30 users in environment 1 repeated the gestures 20 times within 5 seconds, and the remaining 9 users repeated the gestures 20 times within 5 seconds in the other three environments, totaling 6,240 samples.The detailed descriptions of XRF55 dataset are provided in Tab. 5
Appendix B Additional Experiments
Tab. 6 shows that Wi-CBR is comparable to the state-of-the-art competitors within the domain and much ahead in the cross-domain scenarios. Wi-CBR achieves the best results within the domain and in various cross-domain scenarios, even across different datasets. The reason for this is that hand-crafted features are very clever. Still, they may not be able to cover all the domain-independent sequential correlations scattered across the spatio-temporal dimension. For example, the BVP feature is cleverly designed but ignores which cues are essential and which are not. Therefore, we introduce phase and frequency domain data DFS guides containing all the information to capture the spatio-temporal cues in each modality. We also optimize the combined features for cross-fusion to obtain a robust domain-independent characterization.
Table 6: THE accuracy of Wi-CBR under In-Domain, Cross-location, Cross-orientation, and Cross-environment settings in the WIDAR3.0 dataset.
Methods In-Domain CL CO CE
Widar3.0 [7] 92.70% 89.70% 82.60% 92.40%
WiHF with HuFuM [8] 93.92% 90.32% 79.14% 89.67%
WiHF with HuFu [8] 97.65% 92.07% 82.38% --
EI [49] 97.40% 73.33% 79.70% --
WiGRUNT [37] 99.60% 96.49% 93.36% 93.73%
Wi-CBR 99.54% 98.34% 96.30% 96.57%
We evaluate Wi-CBR performance with different users(the default number is 9). The evaluation results are shown in Tab. 7. As the number of users increases, it can be observed that Wi-CBR has an upward trend, and the performance is higher for 9 types of users than for 6 types. This indicates that Wi-CBR can adapt to new users, while we observe that the overall trend of WiGRUNT is decreasing. The performance decreases significantly when the first user is added. We compare the effect of increasing the number of users on the system’s performance with WiGRUNT, and the results are shown in Fig. 5(a), which indicates that Wi-CBR is positively affected by the increase in the number of users.
Table 7: Comparison of Wi-CBR and WiGRUNT performance for different users.
No. of Users Wi-CBR WiGRUNT
6 7 8 9 6 7 8 9
In-Domain 99.07% 99.54% 99.63% 99.54% 99.60% 99.45% 99.48% 99.10%
CL 97.20% 98.46% 98.92% 98.34% 96.49% 96.21% 96.72% 96.17%
CO 95.31% 96.84% 97.08% 96.57% 93.36% 92.86% 94.11% 93.95%
Figure 5: Overall comparison of users and gestures.
We evaluate Wi-CBR performance with different users(the default number is 6). The evaluation results are shown in Tab. 8. Unlike the effect of increasing users, a decreasing trend in performance can be observed for both Wi-CBR and WiGRUNT as the number of gestures increases. However, for intra-domain and cross-location, Wi-CBR decreases slower than WiGRUNT and rises when encountering easily recognizable gestures. For cross-orientation, the two show parallel trends. WiGRUNT continues to decrease throughout, and the results are shown in the Fig. 5(b), which suggests that Wi-CBR’s performance feats are more stable as the number of gestures increases.
Table 8: Comparison of Wi-CBR and WiGRUNT performance for different gestures.
No. of Gestures Wi-CBR WiGRUNT
6 7 8 9 6 7 8 9
In-Domain 99.07% 99.31% 98.67% 98.61% 99.60% 99.26% 98.32% 98.08%
CL 97.20% 97.52% 95.77% 96.03% 96.49% 95.45% 92.25% 93.58%
CO 95.31% 94.89% 91.20% 91.11% 93.36% 92.12% 89.22% 89.01%
We plot cross-domain comparison histograms between Wi-CBR and state-of-the-art methods on the two datasets. The results are shown in Fig. 6, where Wi-CBR leads all previous methods that use unimodal data processing. Our two-branch extraction of spatio-temporal cues module ensures the effective utilization of phase and DFS data. At the same time, cross-modal cross-fusion yields unprecedentedly powerful domain-independent features. A significant advantage is achieved against model-based methods while outperforming the state-of-the-art method Wi-CBR by 8% under the XRF55 dataset.
Figure 6: THE accuracy of Wi-CBR under CL(Cross-location), CO(Cross-orientation), and CE(Cross-environment) settings in the WIDAR3.0 dataset and the XRF55 dataset.

### v4

@@ B.1 Gesture Recognition Dataset Details
Widar3.0 is one of the most widely used datasets for WiFi-based gesture recognition. The dataset collection device consists of one transmitter and six receivers, each equipped with an Intel 5300 wireless card. Each receiver has three antennas arranged in a line. The operating frequency of the device is 5.825 GHz, with a sampling rate of 1000 packets per second. Widar3.0 mainly consists of two types of datasets. Firstly, there are 12,000 gesture samples commonly used in human-computer interaction (16 users × 5 locations × 5 orientations × 6 gestures × 5 instances). Secondly, there are 5,000 gesture samples representing digits 0-9 on a horizontal plane (2 users × 5 locations × 5 orientations × 10 gestures × 5 instances). To ensure a fair comparison, we only utilized 4,500 gesture samples (9 users × 5 locations × 5 orientations × 6 gestures × 5 instances). The detailed descriptions of Widar3.0 dataset are provided in Tab. 9.
XRF55 is a large multi-radio frequency dataset forhuman action analysis. To verify the effectiveness of the proposed Wi-CBR, we selected only the samples related to gesture actions collected using WiFi devices (e.g., drawing a circle, drawing a cross, pushing, pulling, swiping left, swiping right,swiping up, and swiping down) for the experiments. Among them, 30 users in environment 1 repeated the gestures 20 times within 5 seconds, and the remaining 9 users repeated the gestures 20 times within 5 seconds in the other three environments, totaling 6,240 samples.The detailed descriptions of XRF55 dataset are provided in Tab. 10.
B.2 Activity Recognition Dataset Details
XRF55 includes 15 human-object interaction actions, 7 human-human interaction actions, 8 fitness actions, 14 body motion actions, and 11 human-computer instruction actions. These actions were meticulously selected from 19 RF sensing papers and 16 video action recognition datasets. Each action is chosen to support various applications of high practical value, such as elderly fall detection, fatigue monitoring, domestic violence detection,and even accommodating scenarios where one’s hands might be preoccupied or unavailable, like individuals with certain disabilities or temporary injuries. These actions are likely to attract substantial attention from the industrial and academic sectors of the RF sensing community.The detailed activities are shown in Tab. 11.
@@ B.2 Activity Recognition Dataset Details
1st (Classroom) 9
1: Push-Pull; 2: Sweep; 3: Clap; 4: Slide;
5: Draw-O(Horizontal); 6: Draw-Zigzag(Horizontal);
7: Draw-N(Horizontal); 8: Draw-Triangle(Horizontal);
9: Draw-Rectangle(Horizontal)
5 5 10,125
2nd (Hall) 4
1: Push-Pull; 2: Sweep; 3: Clap; 4: Slide;
5: Draw-O(Horizontal); 6: Draw-Zigzag(Horizontal)
5 5 3,000
3rd (Office) 4
1: Push-Pull; 2: Sweep; 3: Clap; 4: Slide;
5: Draw-O(Horizontal); 6: Draw-Zigzag(Horizontal)
5 5 3,000
Table 9: Detailed gesture descriptions of Widar3.0 Dataset
@@
Scene 1 30
1: Drawing circle; 2: Drawing cross;
3: Pushing; 4: Pulling; 5: Swiping left;
6: Swiping right; 7: Swiping up; 8: Swiping down
Changeable Changeable 33,000
Scene 2 3
1: Drawing circle; 2: Drawing cross;
3: Pushing; 4: Pulling; 5: Swiping left;
6: Swiping right; 7: Swiping up; 8: Swiping down
Changeable Changeable 3,300
Scene 3 3
1: Drawing circle; 2: Drawing cross;
3: Pushing; 4: Pulling; 5: Swiping left;
6: Swiping right; 7: Swiping up; 8: Swiping down
Changeable Changeable 3,300
Scene 4 3
1: Drawing circle; 2: Drawing cross;
3: Pushing; 4: Pulling; 5: Swiping left;
6: Swiping right; 7: Swiping up; 8: Swiping down
Changeable Changeable 3,300
Table 10: Detailed gesture descriptions of XRF55 Dataset
1. Shaking hands 20. Playing Ukulele 39. Standing up
2. Hugging 21. Playing drum 40. Sitting down
3. Handing something to someone 22. Jumping 41. Blowing dry hair
4. Hitting someone with something 23. Running 42. Cutting something
5. Choking someone’s neck 24. Turning 43. Drinking
6. Pushing someone 25. Walking 44. Putting something on the table
7. Kicking someone 26. Patting on the shoulder 45. Eating
8. Mopping the floor 27. Foot stamping 46. Carrying weight
9. Combing hair 28. Shaking head 47. Brushing teeth
10. Weightlifting 29. Nodding 48. Using a phone
11. Throwing something 30. Drawing a circle 49. Hula hooping
12. Picking something 31. Drawing a cross 50. Jumping rope
13. Body weight squats 32. Pushing 51. Falling on the floor
14. Waving 33. Pulling 52. Putting on clothing
15. Boxing 34. Swiping left 53. Tai Chi
16. Jumping jack 35. Swiping right 54. Stretching
17. Clapping hands 36. Swiping up 55. Smoking
18. High leg lifting 37. Swiping down
19. Playing Er-Hu 38. Wearing a hat
Table 11: Detailed behavior descriptions of XRF55 Dataset